Electric potential decreases uniformly from 180V to 20 Volt from x = -2 to x = 2. What is the electric field at origin?


 

Grade:12

1 Answers

vijesh jayan
12 years ago

electric field = potential gradient

E=dV/dL

dV is change in potential = 180-20 = 160 V

dL is change in lenght = 2 - (-2) = 4 units

therefore E = 160/4 = 40 V/unit